I cannot understand how to find b in y = mx + b?
as follows:
Question: Put in y = mx + b form, then graph. x-intercept=4, slope=-3
so I wrote it as follows:
y=-3x+b
= 0=-3(4)+b
and here I have the problem, to find b I dont know i I should multiply 3(4) first, or try 0+3=(4)+b... or 0/4=3+b...
I usually come up with b=12, b=-12, b=3/4.... and they say the answer is 4, as in y=-3x+4
Thank you in advance for any aid that you can offer me...
yours truly, Michael

Answer by funmath(2933)

Hi Michael,
If the x-intercept is 4, you're right, b=12. If the y-intercept is 4, then b=4. Check the problem to make sure you copied it right. There could be a type-o in the book.
find b in y = mx + b?
as follows:
Question: Put in y = mx + b form, then graph. x-intercept=4, slope=-3
y=-3x+b
0=-3(4)+b
0=-12+b
12+0=12-12+b
12=b
y=-3x+12
OR if the question was:
Put in y=mx+b form, then graph. y-intercept=4, slope=-3
y=-3x+4
If you need help with the graph, let me know.
Happy Calculating!!!!